OpenStudy (anonymous):

What is the volume of a sphere with a surface area of 9pi yd^2? A) 9pi yd^3 B) 36pi yd^3 C) 9/2pi yd^3 D) 9/4pi yd^3

OpenStudy (anonymous):

C) Solve 4 pi r^2 = 9 pi for r, r = 3/2 v = 4/3 pi r^3 = 4/3 pi (3/2)^3 = (9 pi)/2